Harnett Corporation has two manufacturing departments--Molding and Assembly. The company used the following data at the beginning of the period to calculate predetermined overhead rates:

Molding Assembly Total
Estimated total machine-hours (MHs) 3,000 7,000 10,000
Estimated total fixed manufacturing overhead cost $ 24,000 $ 53,200 $ 77,200
Estimated variable manufacturing overhead cost per MH $ 1.00 $ 2.00

During the period, the company started and completed two jobs--Job E and Job M. Data concerning those two jobs follow:

Job E Job M
Direct materials $ 21,600 $ 9,300
Direct labor cost $ 22,600 $ 9,500
Molding machine-hours 2,500 500
Assembly machine-hours 2,500 4,500

Required:

a. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate that overhead rate. (Round your answer to 2 decimal places.)

b. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate the amount of manufacturing overhead applied to Job E. (Do not round intermediate calculations.)

c. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate the total manufacturing cost assigned to Job E. (Do not round intermediate calculations.)

a. Predetermined overhead rate b. Manufacturing overhead applied c. Manufacturing cost per MH
Answer #1

Answers

  • All working forms part of the answer
  • Working

Working

Molding

Assembly

Total

A

Estimated machine hours

                      3,000

                     7,000

                  10,000

B

Variable overhead per MH

$1

$2

C = A x B

Total Variable Overhead

$3,000

$14,000

$17,000

  • Requirement ‘a’

A

Total variable Overhead

$17,000

B

Tota Fixed Overhead

$77,200

C = A+B

Total manufacturing overhead

$94,200

D

Total machine hours

                    10,000

E = C/D

Plantwide manufacturing overhead rate

$                     9.42 = Answer

  • Requirement ‘b’

A

Job E Total machine hours

$5,000

B

Plantwide manufacturing overhead rate

$                     9.42

C = A x B

Manufacturing Overhead applied to Job E

$47,100 = Answer

  • Requirement ‘c’

A

Direct Materials

$21,600

B

Direct Labor Cost

$22,600

C

Manufacturing Overhead

$47,100

D = A+B+C

Total manufacturing cost assigned to Job E

$91,300 = Answer